OpenStudy (anonymous):

-(3x-2) = -4 (x+1)-4

jigglypuff314 (jigglypuff314):

to solve for x -(3x - 2) = -4(x + 1) - 4 first distribute -3x + 2 = -4(x + 1) - 4 and distribute again -3x + 2 = -4x - 4 - 4 then add like terms -3x + 2 = -4x - 8 then you can add 4x to both sides +4x +4x x + 2 = -8 then subtract 2 from both sides -2 -2 so x = -8 - 2 = (simplify)

OpenStudy (anonymous):

-10?

jigglypuff314 (jigglypuff314):

yep :)

OpenStudy (anonymous):

Thank you

jigglypuff314 (jigglypuff314):

you're welcome :)